OpenStudy (zale101):

Calculus 2: I'm helping someone out regard understanding Maclaurin series. However, they emailed me a question and the wording sort of confused me.

ganeshie8 (ganeshie8):

this is a fun problem recall your old friend, the geometric series : \[\dfrac{1}{1-x} = 1+x+x^2+\cdots\]

OpenStudy (phi):

the first step is replace x with (-u) you get 1-u+u^2-u^3+u^4 ... then replace u with x^2 1-x^2 + x^4 -x^6 etc

OpenStudy (zale101):

I used \(\large f(x)=\frac{a}{1-r}\) \(\large f(x)=\frac{1}{1+x^2}=\frac{1}{1-(-x^2)}\) r=-x^2 a=1

ganeshie8 (ganeshie8):

that looks good to me !

OpenStudy (zale101):

\[\sum_{n=0}^{\ \infty}(-x^2)^n=\sum_{n=0}^{\ \infty}(-1)^n(x)^{2n}=1-x^2+x^4-x^6+..\] That's the expanded version 1/(1+x^2) For clarification, the question wants me to determine the r for 1/(1+x^2) and the replace it with the r in 1/(1-x)?

ganeshie8 (ganeshie8):

yea question is just guiding you to approach the problem in that specific way

OpenStudy (phi):

I interpret it to be asking: Given: 1/(1-x) = 1+x+x^2 + x^3 +... what is the series for 1/(1+x^2) ? answer: rename x as (-x^2) in the given series expansion. We get 1/(1-(-x^2) = 1/(1+x^2) = series

OpenStudy (zale101):

Alright! Made a lot of sense now. Thank you @ganeshie8 and @phi